OpenStudy (anonymous):

Given the following endpoints (A & B) and midpoint (M) on a number line, find the coordinates for A & B. M(-11) A(2x + 3) B(5x + 3) 1) A(-5) & B(-17) 2) A(-17) & B(-5) 3) A(-13) & B(-9) 4) A(-9) & B(-13)

OpenStudy (anonymous):

The distance from A to M is|-11 - (2x + 3)| = |2x + 14|, as is the distance from M to B. If A < M < B, then the coordinate for B is -11 + |2x + 14|. If 2x + 14 <=0, then |2x + 14| = -(2x + 14), and B has coordinate -11- (2x+14) = -25 - 2x. If 2x + 14 >= 0, then B has coordinate -11 + (2x+ 14) = 2x + 3. So, the coordinate for B is -2x - 25 for x<= -7 and 2x + 3 for x >= 7.

OpenStudy (anonymous):

-___- ^^ sooo whats the answer.

OpenStudy (anonymous):

thats what im sayn -_-